twice lisa's age is 11 more than jessica's age. the sum of lisa's age is twice jessica's age is 43. how old are the girls?
Answer by mananth(16949)

let jessica be x
lisa be y
2y=x+11
2y-x=11....................1
y+2x= 43...................2
...
multiply (1) by 2 and add it to (2)
4y-2x=22
add
5y=43+22
5y=65
/5
y=13 lisa's age
...
plug value of y in (1)
26-x=11
26-11=x
x= 15 jessica's age
...
m.ananth@hotmail.ca